OpenStudy (anonymous):

x^2+3 what is the domain, range, interval of increase and decrease,maximum and minimum

OpenStudy (anonymous):

domain=all real numbers range=y greater or equal to 3

OpenStudy (anonymous):

it has a min pt

OpenStudy (anonymous):

minimum = f'(0)=0

OpenStudy (anonymous):

so the domain would be \[-\infty,\infty\]

OpenStudy (anonymous):

f'(x)=0 i mean

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

i dont get the minimum or maximum

OpenStudy (anonymous):

are you in calculus?

OpenStudy (anonymous):

no accelerated math 2

OpenStudy (anonymous):

yes but if she's not in calc, then she doesn't know the derivative lol

myininaya (myininaya):

\[f(x)=x^2+3=(x-0)^2+3\] vertex is (0,3) the graph is concave up so the range is from 3 to infinity [3,inf)

OpenStudy (anonymous):

the inteval of increase is x>0 or x<0

OpenStudy (anonymous):

it can also be said that the range is grater than or equal to 3

OpenStudy (anonymous):

what is the maximum and what is the interval of decrease

myininaya (myininaya):

no max interval of decrease (-inf,3)